On a (very) large island, there are over 10,000 coconut trees. In 2019, you find that 5 of these trees produced the following number of coconuts:

2, 3, 4, 5, 6

Compute the standard deviation of coconuts produced in 2019 from this sample of trees. Round your answer to the 4th decimal place
Answer #1
Values ( X ) Σ ( Xi- X̅ )2
2 4
3 1
4 0
5 1
6 4
Total 20 10

Mean X̅ = Σ Xi / n
X̅ = 20 / 5 = 4


Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 10 / 5 -1 ) = 1.5811